OpenStudy (anonymous):

PLEASE SOMEONE Need Help!! No clue- In alg 2 but looks like calc Find the 7th Partial sum of (see post)

OpenStudy (anonymous):

\[\sum_{i=1}^{\infty} 9i-4\]

OpenStudy (anonymous):

Need help

OpenStudy (anonymous):

If you can solve this, can you, but i am leaving so I won't respond. I will give a medal when I wake up

OpenStudy (zzr0ck3r):

(9(1)-4) + (9(2)-4) + (9(3)-4) + (9(4)-4) + (9(5)-4) + (9(6)-4) + (9(7)-4)

OpenStudy (anonymous):

\[\sum_{i=1}^{4}(9i-4)=9\sum_{i=1}^{4}i-\sum_{i=1}^{4}4=90-16=74\]

OpenStudy (zzr0ck3r):

ahh you did the 4th

OpenStudy (anonymous):

Thank You!